# Spin-glass states generated in van der Waals magnet CrGeTe by alkali-ion intercalation

### v1

@@ 3 Conclusion
In summary, we show the unique feature of Na intercalation processes to significantly modify the transport/magnetic properties of vdW magnetic material CGT. We discovered that the intercalated Na-CGT hosts emerging magnetic states, i.e. the spin-glass states that are absent in pristine CGT. Such a spin-glass phase is a reentrant spin-glass system that coexists with the ferromagnetically ordered background. Dynamic susceptibility measurements analysed by multiple models all suggest the nature of cluster spin-glasses where spins are interacting to each other to behave collectively with the distribution of relaxation timescales. This study sheds light on the unique feature of intercalation to generate magnetically frustrated systems by inhomogeneous carrier doping. This new way of tuning vdW materials might be effective beyond the field of magnetism, where inhomogeneity will play an active role for determining material ground states.
@@ 4 Methods
Direct vapor transport method was used to synthesise the single-crystalline flakes of CrGeTe. Pre-mixing with a molar ratio of of high-purity elemental Cr ( in chips), Ge ( in crystals), and Te ( in beads) was carried out and then sealed in a quartz ampule at high vacuum ( Torr). The ampule was loaded in a single-zone furnace with increasing the temperature at a rate of 2 K/min reaching up to 1273 K. The furnace was set to slow cooling with the rate of 5 K/hour down to 673 K in order to obtain a low defect crystal growth. Then, nature cooling process was used with switching off the furnace. From the flux buildup, the single crystal flakes were separated and then stored in an inert atmosphere in a glovebox.
CGT was de-gassed in vacuum to a pressure less than 10 mbar at 100C. Na and CGT were placed in a custom-made air-tight glass tube at C and submerged in liquid ammonia, with a 1:1 molar ratio of Na:CrGeTe. The solution was maintained for a period of approximately 72 hours to allow for the intercalation to occur.
X-ray diffraction was performed using a Philips X’pert MPD diffractometer with Cu K radiation ( Å) in reflection geometry. For the measurement on air-sensitive Na-CGT, the sample was loaded into a berylium dome holder with O-ring type seal inside an Ar glovebox for measurement. Raman spectroscopy was performed using a Renishaw In-Via microscope equipped with a 488 nm laser and a 2400 I/mm grating, with a x20 magnification. Samples were measured in an air-free glass sample holder, loaded in a glovebox.
@@ Acknowledgements
S.K. and H.K. acknowledge support from the UK Engineering and Physical Sciences Research Council (EPSRC) via EP/T006749/1. E.S.Y.A. would like to thank the EPSRC for studentship funding from the Centre for Doctoral Training in Advanced Materials Characterisation (EP/S023259/1). L.A.V.N-C acknowledges a scholarship EP/R513180/1 to pursue doctoral research from EPSRC. A.S. thanks JSPS Postdoctoral fellowship for research in Japan (P21777). Magnetic measurements at the University of Cambridge were made on the EPSRC Advanced Characterization Suite (funded under EP/M0005/24/1). E.J.G.S. acknowledges computational resources through CIRRUS Tier-2 HPC Service (ec131 Cirrus Project) at EPCC (http://www.cirrus.ac.uk) funded by the University of Edinburgh and EPSRC (EP/P020267/1). EJGS acknowledges the Edinburgh-Rice Strategic Collaboration Awards and the EPSRC Open Fellowship (EP/T021578/1) for funding support. A.G.L. acknowledges financial support from the Spanish Ministry of Science, Innovation and Universities under grant No. PRX19/452.
@@
S.K. and H.K. conceived the idea of the project. I.V. and G.E. synthesised and provided pristine CGT crystals. E.S.Y.A. and C.A.H. intercalated Na-ions in the CGT crystals. E.S.Y.A. performed Raman and XRD experiments. L.A.V.N-C. conducted dc-magnetometry and ac susceptibility experiments with supervision of S.E.D. S.K. performed the electrical transport and magnetotransport experiments. Data analysis and interpretations were carried out by S.K., E.S.Y.A., L.A.V.N-C., H.K. and all other co-authors. S.K., E.S.Y.A., L.A.V.N-C., A.S. and H.K. wrote the manuscript with input from the other co-authors.
Supplementary Information for “Spin-glass states generated in van der Waals magnet CrGeTe by alkali-ion intercalation"
@@
In order to distinguish background peaks from CGT 00l diffraction patterns, a graphite reference was also measured in a beryllium dome. The peaks shown in Figure S1 (a) corresponds to graphite peaks from literature, with the exception of the peak at 44 which is present in both diffraction patterns. Hence, the peak at 44 can be attributed to the sample environment, as opposed to Na-CGT.
@@
The analysis of the Raman spectroscopy data was performed using the Python 3 S[5] package SciPy S[6]. Using the curve_fit function, a Lorentzian peak-shape was fitted to each of the modes studied, defined as:
@@
S3 DC-SQUID Measurements
As highlighted in the main text in Figure 2(a), the - curve in Na-CGT shows a different behaviour compared to CGT. In Figure S4, the magnetisation response as a function of temperature is compared along the two orientations of out of plane (c-axis) and in-plane (ab-plane) for the Na-CGT sample. It is evident from this measurement that the magnetic ordering temperature has been enhanced and the value of goes to zero in the vicinity of 240 K (inset of Figure S4).
@@
As mentioned in the main text, the ac magnetic susceptibility is expressed in complex form as , where the real part ,, is the in-phase component and the imaginary part, , is the out of phase component S[10]. Figure S8 shows the temperature dependence of for different excitation frequencies, and even though the amplitude is smaller compared to the real part shown in Fig. 3(b) in the main text, the anomaly feature is clearly present in , supporting the spin-glass picture. In typical spin-glass systems, relative low amplitude of is suggestive of distribution of relaxation mechanism instead of a single relaxation time S[11].
@@
where and is Boltzmann constant. As pointed out in the main text, even though the experimental data can be fit using Eq.(S3), the best fit parameters of = 8035 K and are both unphysical. The relaxation mechanisms in Na-CGT can thus not be described in terms of non-interacting spins, suggesting the interacting nature of Na-CGT’s magnetic clusters S[11].

### v3

@@ 3 Conclusion
In summary, we show the unique feature of Na intercalation processes to significantly modify the transport/magnetic properties of vdW magnetic material CGT. We discovered that the intercalated Na-CGT hosts emerging magnetic states, i.e. the spin-glass states that are absent in pristine CGT. Such a spin-glass phase is a reentrant spin-glass system that coexists with the ferromagnetically ordered background. Dynamic susceptibility measurements analysed by multiple models all suggest the nature of cluster spin-glasses where spins are interacting to each other to behave collectively with the distribution of relaxation timescales. This study sheds light on the unique feature of intercalation to generate magnetically frustrated systems by inhomogeneous carrier doping. The ammonia-solution-based, alkali-atom intercalation techniques presented in this study can be performed with other alkaline elements such as K and Ca, for further exploring the tunability of magnetic ground states and induced frustration with different intercalants. We envisage that unique spin-glass properties generated by this alkali-atom intercalation route will be actively discussed and identified once a good number of experimental demonstrations have been reported. This new way of tuning vdW materials might be effective beyond the field of magnetism, where inhomogeneity will play an active role for determining material ground states.
@@ 4 Methods
Direct vapor transport method was used to synthesise the single-crystalline flakes of Cr2Ge2Te6. Pre-mixing with a molar ratio of of high-purity elemental Cr ( in chips), Ge ( in crystals), and Te ( in beads) was carried out and then sealed in a quartz ampule at high vacuum ( Torr). The ampule was loaded in a single-zone furnace with increasing the temperature at a rate of 2 K/min reaching up to 1273 K. The furnace was set to slow cooling with the rate of 5 K/hour down to 673 K in order to obtain a low defect crystal growth. Then, nature cooling process was used with switching off the furnace. From the flux buildup, the single crystal flakes were separated and then stored in an inert atmosphere in a glovebox.
We used established low-temperature intercalation techniques [71, 46] to synthesise Na-CGT. Macroscopic CGT crystals were outgassed under vacuum to a pressure less than 10-6 mbar at 100∘C overnight, in a custom-made air-tight glass tube. In an Ar glovebox, the mass of outgassed crystal was measured, followed by adding Na into the glass tube at a molar ratio of Na:CGT = 1:1. To be specific, we used mass ratio of Na:CGT = 4.7 mg : 208 mg and another batch of 4.3 mg: 184.3 mg. The glass tube was then submerged in a bath at approximately -50∘C, where liquid ammonia was condensed onto CGT and Na. The solution was maintained for a period of approximately 72 hours to allow the intercalation to fully complete to form a crystal with the atomic ratio of Na1Cr2Ge2Te6. We can confirm the completion of intercalation by the color change of the ammonia solution [46, 72]. During intercalation, the alkali-metal-ammonia solutions display a deep blue color, arising from the solvated electrons. As the reaction progresses the blue colour is gradually lost as the electrons move onto the layers driving concomitant intercalation. Eventually we observed an orange solution, which can be attributed to the presence of sodium polytelluride excess byproduct. Thus, we can confirm that the intercalation of sodium into the CGT matrix has occurred. Subsequently, the ammonia was slowly extracted via cryo-pumping until intercalated Na-CGT crystals were left.
X-ray diffraction was performed using a Philips X’pert MPD diffractometer with Cu Kα radiation ( Å) in reflection geometry. For the measurement on air-sensitive Na-CGT, the sample was loaded into a berylium dome holder with O-ring type seal inside an Ar glovebox for measurement. Raman spectroscopy was performed using a Renishaw In-Via microscope equipped with a 488 nm laser and a 2400 I/mm grating, with a x20 magnification. Samples were measured in an air-free glass sample holder, loaded in a glovebox.
@@
Theoretical modeling of the exchange interactions before and after the Na-intercalation of CGT was undertaken to elucidate the main factors inducing the appearance of competing spin interactions (e.g., isotropic and anisotropic exchange, anisotropy energies, Dzyaloshinskii-Moriya interaction - DMI) responsible for the formation of the spin-glass state, enhancement of , and change of magnetic axis. See Supplementary Information Sec S7 for details.
@@ Acknowledgements
S.K., C.A.H. and H.K. acknowledge support from the UK Engineering and Physical Sciences Research Council (EPSRC) via EP/T006749/1 and EP/V035630/1. E.S.Y.A. would like to thank the EPSRC for studentship funding from the Centre for Doctoral Training in Advanced Materials Characterisation (EP/S023259/1). L.A.V.N-C acknowledges a scholarship EP/R513180/1 to pursue doctoral research from EPSRC. A.S. thanks JSPS Postdoctoral fellowship for research in Japan (P21777). Magnetic measurements at the University of Cambridge were made on the EPSRC Advanced Characterization Suite (funded under EP/M0005/24/1). E.J.G.S. acknowledges computational resources through CIRRUS Tier-2 HPC Service (ec131 Cirrus Project) at EPCC (http://www.cirrus.ac.uk) funded by the University of Edinburgh and EPSRC (EP/P020267/1). EJGS acknowledges the EPSRC Open Fellowship (EP/T021578/1), and DIPC for funding support. A.G.L. acknowledges financial support from the Spanish Ministry of Science, Innovation and Universities under grant No. PRX19/452.
@@
S.K. and H.K. conceived the idea of the project. I.V. and G.E. synthesised and provided pristine CGT crystals. E.S.Y.A. and C.A.H. intercalated Na-ions in the CGT crystals. E.S.Y.A. performed Raman and XRD experiments. E.J.G.S. and S.G. provided theoretical modeling and contributed to the explanation of the spin-glass state. E.S.Y.A. and L.A.V.N-C. conducted EDX characterisation and analysis. L.A.V.N-C. conducted dc-magnetometry and ac susceptibility experiments with supervision of S.E.D. S.K. performed the electrical transport and magnetotransport experiments. Data analysis and interpretations were carried out by S.K., E.S.Y.A., L.A.V.N-C., H.K. and all other co-authors. S.K., E.S.Y.A., L.A.V.N-C., A.S. and H.K. wrote the manuscript with input from the other co-authors.
Supplementary Information for “Spin-glass states generated in a van der Waals magnet by alkali-ion intercalation"
@@
In order to distinguish background peaks from CGT 00l diffraction patterns, a graphite reference was also measured in a beryllium dome. The peaks shown in Figure S1 (a) corresponds to graphite peaks from literature, with the exception of the peak at 44∘ which is present in both diffraction patterns. Hence, the peak at 44∘ can be attributed to the sample environment, as opposed to Na-CGT.
@@
The analysis of the Raman spectroscopy data was performed using the Python 3 S
[5] package SciPy S
[6]. Using the curve_fit function, a Lorentzian peak-shape was fitted to each of the modes studied, defined as:
@@
S3 Energy dispersive X-ray analysis
Electron dispersive x-ray spectroscopy (EDX) was performed to investigate the local stoichiometry of Na-intercalated CGT. A Tescan MIRA3 SEM was used for the measurements. To avoid air exposure and subsequent sample degradation, each Na-intercalated CGT sample was loaded into an airtight sample-holder and transported to the SEM. It was not necessary to coat samples with conductive material and an accelerating voltage of 10 kV was used.
As discussed in Method section in the main manuscript, the expected stoichiometric ratio was 1:2:2:6 for Na:Cr:Ge:Te, from the ratio of source materials used in the Na intercalation. Figure S4(a) shows a representative EDX spectrum at point 21 in the SEM image in Fig. S4(b), which shows the presence of Na (8.5at%), Cr (17.8at%), Ge (17.0at%), and Te (56.7at%), which conforms successful Na intercalation that produces a similar composition ratio to the expected stoichiometry. It is worth noting that the Na peak is relatively weak, impacting on the accuracy of the Na atomic percentage and this ratio. The inherent lack of accuracy of SEM-EDX for elemental analysis is also evident in the Cr:Ge ratio, although such variation is common in SEM-EDX measurements due to a range of factors S [10]. Element-sensitive mapping of the Na-CGT as shown in Fig. S4(c)-(f) show each element is distributed across the sample, with intensity variation typical for SEM-EDX maps, arising from uneven surface topography S [10].
S4 DC-SQUID Measurements
As highlighted in the main text in Fig. 2(a), the - curve in Na-CGT shows a different behaviour compared to CGT. In Figure S5, the magnetisation response as a function of temperature is compared along the two orientations of out of plane (c-axis) and in-plane (ab-plane) for the Na-CGT sample. It is evident from this measurement that the magnetic ordering temperature has been enhanced and the value of goes to zero in the vicinity of 240 K (inset of Fig. S5).
@@
As mentioned in the main text, the ac magnetic susceptibility is expressed in complex form as , where the real part ,, is the in-phase component and the imaginary part, , is the out of phase component S
[11]. Figure S9 shows the temperature dependence of for different excitation frequencies, and even though the amplitude is smaller compared to the real part shown in Fig. 3(b) in the main text, the anomaly feature is clearly present in , supporting the spin-glass picture. In typical spin-glass systems, relative low amplitude of is suggestive of distribution of relaxation mechanism instead of a single relaxation time S
[12].
@@
where and is Boltzmann constant. As pointed out in the main text, even though the experimental data can be fit using Eq.(S2), the best fit parameters of = 8035 K and are both unphysical. The relaxation mechanisms in Na-CGT can thus not be described in terms of non-interacting spins, suggesting the interacting nature of Na-CGT’s magnetic clusters S
[12].
@@
S7 Theoretical Methods
S7.1 Ab initio simulations
To provide additional insights of the spin-glass states in Na-CGT systems, we undertook multi-scale simulations using density functional theory (DFT) to calculate the magnetic parameters (e.g., exchange interactions, anisotropies, etc.), and Monte Carlo to simulate variations of Curie temperature. On the DFT part, geometry optimizations of bulk Cr2Ge2Te6 are performed using Vienna Ab initio Simulation Package (VASP)S [14, 15]. The Na intercalation is performed by changing the position of Na in the unit cell and the magnetic parameters are evaluated for the lowest energy configuration. The generalized gradient approximation (GGA) has been used to treat the exchange-correlation interactions within the Perdew-Burke-Ernzerhof (PBE) formS [16]. We use 18186 Monkhorst-Pack -point mesh in our calculations for Brillouin zone (BZ) integration.S [17] The lattice parameters and atomic coordinates are optimized by the energy minimization technique based on the conjugate gradient algorithm with a force component tolerance of 0.01 eV/Å on each atom. The cutoff for the energy of plane-wave basis set is considered to be 500 eV.
The full-potential linear muffin-tin orbital (FP-LMTO) method, implemented in the RSPt codeS [18] is used to study the magnetic and electronic properties of Fe5-δNiδGeTe2 monolayers. First we perform standard density functional theory (DFT) calculations, The converged DFT calculations are the starting point to perform our DFT+U calculations as implemented in RSPtS [18]. The value of Hubbard parameter or ( is the Hund’s exchange) for Cr-3 states is used to be 0.5 eV, similar with the previous studiesS [19].
S7.2 Calculation of exchange interactions
The isotropic symmetric exchange interactions are calculated within the full-potential linearized muffin-tin orbital (FPLMTO) basis implemented in the RSPt code. From the LMTO basis, one can construct the Bloch sums to solve the DFT eigenvalue problem and subsequently for the one-electron Green’s function.
The couplings are computed using force-theorem based Green’s function formalismS [20]. The intersite exchange parameters are calculated using eq. S8:
where and are the onsite spin splitting and the spin-dependent intersite Green’s function, respectively. The trace in eq. S8 is over the orbital degrees of freedom. and are the th fermionic Matsubara frequency and the temperature, respectively. The onsite exchange splitting term includes the self-energy, which is given by:
where, and are the Kohn-Sham Hamiltonian and site-dependent self-energy. The self-energy is obtained by solving the DMFT equations. In DMFT calculations, The frequency-dependent self-energy is obtained from DMFT. The exchange splitting also depends on frequencyS [20].
In the fully-relativistic case the generalized Heisenberg model becomes:
where , . The magnetic exchange parameters are () tensors in the considered fully-relativistic case, which is a generalization of the scalar parametersS [20]. To calculate the magnetic exchange interactions we use k-point mesh.
S7.3 Exchange interactions in Na-intercalated CrGeTe3 magnets
In the following we calculated the isotropic exchange interactions (Table S2), antisymmetric exchange interactions (e.g., Dzyaloshinskii-Moriya - ) (Table S3), and anisotropic exchange (Table S4). The inclusion of Na dopants break the symmetry of the Cr atoms, which generated additional interactions responsible for likely competition between pristine areas where ferromagnetism is present, and areas where Na-induced potential frustrations. Figure S13 shows a schematic of the CGT layers with intercalated Na atoms into the system.
| System | ||||||||
| Pristine | 5.478 | -0.019 | 0.687 | -0.052 | 0.121 | 0.302 | -0.068 | -0.053 |
| Na-intercalated | 5.325 | 0.790 | 1.667 | 0.227 | -0.162 | 0.471 | 0.259 | 0.4689 |
| 4.369 | -0.171 | 1.024 | -0.297 |
| System | ||||||||
| Pristine | 0.000 | 0.159 | 0.000 | 0.000 | 0.000 | 0.005 | 0.003 | 0.000 |
| Na-intercalated | 0.309 | 0.165 | 0.149 | 0.028 | 0.177 | 0.000 | 0.128 | 0.102 |
| 0.000 | 0.186 | 0.000 | 0.020 |
| System | ||||||||
| Pristine | 0.192 | 0.003 | 0.036 | 0.008 | 0.000 | 0.008 | 0.005 | 0.002 |
| Na-intercalated | 0.129 | 0.017 | 0.049 | 0.003 | 0.000 | 0.001 | 0.007 | 0.006 |
| 0.156 | 0.0178 | 0.0795 | 0.005 |
S7.4 Magnetic Anisotropy
The single-ion anisotropy is estimated based on the one-shot fully-relativistic calculations for the spin axis pointing along the , , and directions. These calculations are run starting from the fully-converged self-consistent non-relativistic electronic structure. The value of anisotropy is obtained from the difference in the sum of the energy eigenvalues for the three spin directions, as already mentioned. Note that, this approach, which is based on the force theorem, often is a more accurate way of determining the magnetic anisotropy, compared to relativistic total energiesS [21]. We use the k-point grids with dimensions for the magnetic anisotropy calculations. A change in orientation from easy-axis to easy-plane is observed after the introduction of Na atoms.
| System | K |
| Pristine | -0.08 |
| Na-intercalated | 0.23 |
S7.5 Monte Carlo simulations
Solving the Heisenberg spin Hamiltonian, considering localized spin moments, the Curie temperature is computed performing Monte Carlo simulations:
where , and are isotropic symmetric, antisymmetric and anisotropic symmetric interactions between th and th species, is the single-ion anisotropy. Such Hamiltonian has been intensively investigatedS [22, 23, 24, 25, 26, 27, 28, 29, 30, 31] throughout many systems and conditions which provide a robust background for the simulations.
To estimate the Curie temperatures () shown in Figure S14, we performed classical Monte Carlo (MC) simulations via UppASD codeS [32], where the calculated magnetic parameters are implemented in Eq. S11. To achieve properly averaged properties, calculations are done for three ensembles in the supercell with size 303010, where periodic boundary conditions are imposed along and axes. is estimated by monitoring the cross sections of fourth-order cumulants of magnetization. The number of Monte Carlo steps considered in our calculations is .